OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

@sourwing @satellite73 @shamil98

OpenStudy (anonymous):

\[7x^2-14x+7=0\\ x^2-2x+1=0\\ (x-1)^2=0\\ x=1\]

OpenStudy (anonymous):

so in your case \[\sin(x)=1\]

OpenStudy (anonymous):

Is that it? @satellite73

OpenStudy (anonymous):

what is what?

OpenStudy (anonymous):

I'm trying to do trigonometric equations. Is pi involved here?

OpenStudy (anonymous):

do you know a number \(\theta\) with \(\sin(\theta)=1\)? there are infinite such numbers

OpenStudy (anonymous):

Im confused. I know that symbol is theta

OpenStudy (anonymous):

forget the symbol solve for \(x\) \[\sin(x)=1\]

OpenStudy (anonymous):

x=1?

OpenStudy (anonymous):

Im lost :(